# Rui Eduardo Lopes, Student Member, IEEE, Susana Sargento, Senior Member, IEEE

### v1

@@ Abstract
With modern vehicles evolving with more features, services, complex systems, with more sensors, actuators, and processing units, it is essential to think about vehicles not only as means of transportation that may tend towards full autonomy, but also as adaptive objects, that suit themselves to the needs of occupants. Vehicular services can be developed to support these adaptations. However, the increasing complexity of vehicular service development, even with current standardizations and best practices and guidelines, are insufficient to tackle the high complexity of development, with expectations of up to 1 (U.S.) billion lines of code for a fully (level 5) autonomous vehicle.
Within this survey, the paradigm of Deterministic Software Defined Vehicles is explored towards increasing the quality and easiness of the development of services for automotive.
Towards this, a proposed vision with four pillars is also provided: the deterministic network configurator, the data layer configurator, and the hypervisor configurator and the vehicle abstraction layer, all coordinated by a software orchestrator.
Index Terms:
Software-Defined Vehicles, Connected Vehicles, Vehicular Applications, Vehicular Services, Intelligent Transportation Systems, Deterministic Networks, Time Sensitive Networks
@@
I Introduction
We are living in an ever more connected world, which is not very safe nor designed for everyone. This is particularly true within vehicular scenarios, where autonomous next generation vehicles are designed solely to remove the need of a driver without much consideration for occupants and other road users’ specific needs. This is even more substantial in a more and more elderly society, where impairments are unavoidable.
Even if vehicle will tend to become driver-less, autonomous vehicles may also take decades to be fully on the roads and starting now, they will coexist with manual-driven cars. As such, autonomous vehicles do not take on the needs of the other occupants. Examples of such needs would be the vehicle recognizing them or potential health issues and providing personalized profiles and automatic adjustments on seating, cooling systems, driving profiles but also other vehicular subsystems that may support some sort of impairment.
It is therefore essential to think about vehicles not only as means of transportation that may tend towards not requiring drivers, but also as adaptive objects, that suit themselves to the needs of occupants. But this adaptation must be carried out flawlessly, since any failure can mean harm to someone, or even death. The context of deterministic — i.e. where the result is exactly what was intended in the exact moment it was intended — critical vehicular services — pieces of software that carry out something within the vehicle or its environment — within Software-Defined Vehicle can provide a solution to this problem.
@@ I Introduction
With the introduction of computational systems into vehicles in late 1980s, the complexity — i.e. , the amount of systems that need to be choreographed, developed, tested and that have to communicate — of vehicles exploded, as well as the possibilities of something going wrong [2, 3]. Product compliance in the automotive industry already have to deal with more than 100 million source lines of code [4] and will have to deal with more than 300 million lines of code by 2030, to take into account more than 500 thousand regulations and requirements (80% of them focused on safety) [5]. With fully autonomous driver-less cards, Original Equipment Manufacturers estimations are even more problematic, with expectations of up to 1 billion lines of code [4].
Current vehicular systems are extremely complex, with many Electronic Control Units (at least around 100111Described in www.fortunebusinessinsights.com.), sensors, and actuators scattered throughout the vehicle and more than 100 million lines of code [^2] More than in a military or commercial aircraft (F-35, 787 Dreamliner)  [6]., with perspective of reaching 500 million to 1 billion with level 5 autonomous vehicles [6]. This is part of an ever-growing tendency of increasing complexity in vehicular architectures, as depicted in fig. 1.
This growing code base is becoming unmanageable to vehicle manufacturers and developers of vehicular services. As described in fig. 2, automakers will tend to leave unchecked, unverified and poorly tested software go into production vehicles if the growth trend of the code base continues — with potentially deadly results [^3] Unverified logic and deadly effects on Tesla’s autopilot www.washingtonpost.com.. The overall intra-vehicular systems as developed following legacy standards and practices are deemed too complex and static — and thus costly to maintain, update and keep safe — by vehicle manufacturers, with the reduction of the overall complexity (both in software and hardware terms, in particular wiring harness weight savings) becoming a major goal in order to reduce operational costs and decrease potential security issues from unchecked, poorly tested code [7].
Standardization of best practices and development guidelines for intra vehicular software gained traction during the early 2000s [8], they are insufficient to tackle such complexity of development. While programming methodologies [^4] Such as static code analysis, Test Driven Development (TDD) and equivalent methodologies. and guidelines [^5] Such as the Automotive Safety Integrity Levels within ISO 26262 or Motor Industry Software Reliability Association (MISRA)-C, a set of guidelines for safety-related systems, such as automotive vehicles. Details at synopsys.com, misra.org.uk and synopsys.com. are critical and cannot be disregarded, they are no longer enough to tackle this complexity.
In this article, we propose that the quality and easiness of the development of services for automotive can be improved using the paradigm of Deterministic Software Defined Vehicles. The main contributions of this tutorial are as follows:
@@
    Demonstrate the suitability and relevance of the software defined vehicles for an sustainable development of vehicular services.
@@
The remaining of this document is organized into four sections. Section II: Software Defined Vehicles — introduces the background concepts, defining vehicular services and their characteristics and the concept of Software-Defined Vehicle (SDV) and additional information on the industrial landscape and how automakers or suppliers are facing the SDV paradigm. Next, section III: Main building blocks of Software Defined Vehicles — presents the main building blocks for a SDV based vehicle (and vehicular services) development, and section IV: Open challenges — discusses on the main challenges.
Section V: Potential use cases — follows with the main use cases and requirements that can be drawn from them, while Section VI: Proposed vision — considers those requirements for a proposed architecture, as well as a proposal of how the evaluation process to validate the work may be considered, with considerations on required hardware and software for results obtained both through simulation or real-world testing. Finally, conclusions are provided in section VII: Conclusions.
@@
II Software Defined Vehicles
This section presents the concept of vehicular services, software defined vehicles and the efforts in the framework of the industry landscape.
II-A Vehicular Services
The desired transition from fossil-fueled vehicles towards clean-energy Intelligent Transport System (ITS) requires the introduction of more advanced systems and services. Such services are required to support the necessary auxiliary electrical systems, increase overall range while decreasing the environmental impact, but also to improve the overall quality of life for passengers and drivers and support (semi)autonomous driving. Vehicular services can be drawn a parallelism to personal computing applications: some services are critical, others are less critical, some require vast computing resources, either in terms of Central Processing Unit (CPU), memory, storage, or graphical resources [10], and can broadly be classified in terms of real-time guarantee levels in [11] either hard real-time services (where deadline failure means a system failure with potential catastrophic consequences), weakly hard real-time (where some tasks can be skipped, while keeping the deadline fulfillment) or soft real-time (where a specific number of deadlines can be missed).
Critical services such as safety services (e.g. emergency braking) can be considered within the SDV as well as services related to the occupants [12]. Main requirements of most of those services include high interoperability, reliability, availability, safety, security, performance, scalability, extensibility, adaptability and testability [13]. During the overall definition and characterization of a vehicular service it is essential to formulate the service and characterize its tasks in terms of those requirements, with particular emphasis on characteristics such as the priority of the service, data specifications — such as format, size and real time guarantee requirements (periodicity or deadline for receiving/sending data) — expected computing location (vehicle edge vs smart city cloud) [11, 14]. Non functional characteristics such as cost minimization, efficient use of available resources, maintainability and scalability, high flexibility to add new services (without custom hardware and software stacks), and redundancy at network and computing level, while keeping backwards compatibility with older vehicles shall also be taken into account [14].
These characteristics can be seen as the Service Level Agreement (SLA) of the service. Such SLAs can and should be defined by the developer of each vehicular service and could be ensured/enforced by a distributed configurator through a set of Quality of Service (QoS) tools and policies, including [11]:
  • Network and resource monitoring and discovery.
  • Dynamic, intelligent task scheduling and resource estimation, provisioning and allocation, including the offloading to other layers of communication (fog and cloud), if the overall end to end communication latency is capable of supporting such offloading while fulfilling deterministic well defined deadlines.
  • Federation of multiple physical wireless and wired access layers in order to obtain a multi-homing solution.
  • Slicing of the multi-homing communication solution to support different levels of QoS specified by the service descriptors.
  • Management of the high mobility, for example through the management of reactive and proactive handovers through Software-Defined Network (SDN[15] based solutions.
  • Prediction of the overall resource usage while considering heterogeneity of the nodes. For example, a system that, through the service descriptor and usage trends of eacfh/all ECUs, analyzes the likelihood of such service of being resource starved or not, can be a good first estimation [^6] Author proposal.. This can be further improved by defining an overall cost of each node [16], or by introducing reinforcement learning techniques into the scheduling of the tasks [17].
  • Estimation and prediction of the mobility of a certain service deployment based on trends of resource usage and prediction of the mobility of the vehicle [18].
While considering specifics of the vehicular domain, such as the mobility of vehicular nodes, communication and node failures, heterogeneity of the nodes hardware specifications, overall low specification and communication techniques and overall security threats.
II-B Concept
@@
SDV aims to improve cars along their life-cycle [19], while potentially mitigating the increased unmanaged complexity of modern vehicular software, hardware, and communications. The paradigm states that vehicles manage their operations, add functionality, and enable new features primarily or entirely through software [^7] More details at https://blackberry.qnx.com/en/ultimate-guides/software-defined-vehicle.. Since vehicles evolve over their lifetime, with more features, solutions that may require remote control or remote updates, main features being delivered by software adds flexibility and reduces the overhead.The shift towards SDV signifies new features can be pushed to existing and new vehicles, which are now more defined through software than through hardware or mechanical systems as in the past [6]. To support this new paradigm, computational architectures within vehicles are going from fully distributed ECUs scattered throughout the vehicle towards a more centralized solution, with zone-based architectures, i.e. ECUs positioned in specific parts of the vehicle responsible for certain types of actions [20]. ECUs are now responsible for more functions, thus being more powerful, powered with tailor made System-on-chips with more cores and Digital Signal Processors, Graphical Processing Units and Ethernet capability [6] with partnerships with Integrated Circuits manufacturers [21].
@@
SDV support not just the improvement of the development of vehicular services but also helps towards a decrease in hardware complexity (with decreased number of ECUs, cables and the cost and weight savings which follow), with more uniform software code-base (and the reduction of the costs and complexity of the development process which follow) and greater adaptability [^8] This adaptability is essential specially if the future of the car ownership is car sharing [22]. towards the future, with the possibility of new or updated features as over the air updates during the vehicle’s lifetime. This adaptability also supports a greater capability to process higher volumes of data while fulfilling the overall performance requirements — essential in Connected Autonomous Vehicles.
Vehicles in the past added new features at the cost of new specialized ECUs close to the sensors and actuators and connected to them through CAN or other equivalent bus (more details at section III-D. This tendency created an upwards trend in the number of ECUs and connecting wiring.
The new Electrical/Electronic (E/E) architectures within a vehicle follow the principle of aggregation of ECUs into a small subset of more powerful ECU, interconnected through a high performance bus, such as Ethernet [23]. Figure 3 provides an overview of this evolution.
Figure 3: Legacy, current, and future SDV paradigm on the topology of ECUs within a vehicle.
Intra vehicular computational architectures have evolved from the original distributed approach, where custom isolated ECUs implement a dedicated very specific function towards interconnected ECUs with various degrees of aggregation of function. With the introduction of ECUs during the 1980s, as in fig. 3 (a), ECUs did very specific functions in an isolated manner and connected only to specific sensors and actuators through Controller Area Network bus (CAN) or equivalent. With the explosion of features in vehicles, specially luxury cars, during the 1990s, ECUs became interconnected. By 2000s, the high complexity of intra vehicular systems became subject to de facto standards such as Automotive Open System Architecture (AUTOSAR[2, 3] and buses such as FlexRay aimed at mitigating some of the shortcomings of CAN.
Decentralized architectures were the norm within intra vehicular systems, allowing a better decoupling, separation of concerns and higher tolerance to equipment failure. However, the scalability of the deployment of services to multiple ECUs (specially considering all the possible ECU variations), the performance of communications and the overall cost in bulky wiring harnesses and several ECUs created a trend towards a more centralized architecture, with more centralization of processing in less computing units. This follows previous trends within other industries, in particular the aviation industry [14].
While distributed architectures tend to ensure high reliability in the event of equipment failure and each ECU is simpler, cheaper, and with a very specific purpose (and thus easier to outsource its manufacturing), the overall architecture is not designed for scalability, flexibility and changeability. By 2010s, domain architectures were a first step towards mitigating those issues through the aggregation of functions per domain. However, the integration of multiple domains required by systems such as Advanced Driving Assistance System (ADAS) require coordination between different domains [2, 3].
A fully centralized architecture aims at mitigating that issue while potentiating reduction of costs in the number of ECUs and the higher software reusability, while allowing greater flexibility. In a centralized architecture, high performance microprocessor [^9] Microcontroller based ECUs are only for very specific safety real-time subsystems. based ECUs are interconnected through a high performance bus [2, 3]. Centralization within the intra-vehicular systems can thus be achieved at a domain, cross domain or zone/vehicle centralized way [14]:
  • Domain oriented solution: In a domain-oriented solution, the clustering is done by vehicle function, which requires heavy interaction between ECUs for complex systems such as ADAS.
  • Cross domain solution: In a cross-domain solution, functions of several domains are consolidated within more powerful ECUs. This minimizes, without removing, the high bandwidth required to transfer between ECUs.
  • Zone-oriented architecture: In a zone-oriented architecture, ECUs in each zone of the vehicle collect and forward information to a much more powerful central ECU, or even offloads to the cloud. Bosch reported on achieving a 15–20% reduction in wiring harness weight when using a zone-oriented architecture instead of a domain-oriented architecture.
Since functions are now centralized in less ECUs, each one must be more powerful, with multi core CPUs, GPUs, DSPs, and potentially Field-Programmable Gate Arrays. To ensure optimal usage of ECUs, with concurrent execution of different services with different criticality, partitioning of scheduling of each core per service(s) is required. Virtualization of resources, either memory, buses (e.g. CAN) and CPU allows both isolation of services within virtual machines while allowing migration and reusage of legacy software (by installing legacy stacks on top of the virtual machines) [14]. Some ECUs can also be dedicated to computing intensive tasks, including Machine Learning (ML) and Deep Learning (DL) workloads supporting ADAS systems [13].
II-C Industrial landscape
@@
Vehicular manufacturers are aware of the necessary changes in thinking. The industry, while already aware for the need of such solutions, the transition from the previous mentality of closed source, customized solutions for each model of vehicle towards a more extendable, interoperable, architecture. In particular, the Renault, Volvo/Polestar and Volkswagen group seems to be ahead in terms of a more interoperable architecture based on ECUs interconnected through Ethernet that run Android and well-known versions of chipsets. It is however not possible to ignore efforts from other vehicle manufacturers that may be developing equivalent solutions but have not, at the time of this survey, announced their efforts. The opacity of the vehicular industry — expected within a competitive market — while partially mitigated through the research of gray literature, remains and will remain a threat to any related work, including this work.
After the landscape changing forced by Tesla, automakers such as BMW [24]; the Renault group through the Renault Software Factory [^10] More details on www.reuters.com.; Volvo; Kia and Hyundai [^11] More details on www.carscoops.com.; and the Volkswagen group through the CARIAD [^12] Available at cariad.technology., supported by companies such as Vector [25] and VXworks, are shifting towards a software-first development of the vehicles, based on novel hardware and communication paradigms. The next phases of automotive software development should also have a deep integration with cloud and follow the best practices of software development (CI/CD) [26]. Figure 4 provides an overview of the efforts of automakers on adoption the main aspects related to SDV, with table I providing further detail per associated areas.
@@
TABLE I: Overview of the manufacturers effort on SDV.
@@
The Tesla company is probably the most well-known example of application of the SDV paradigm principles. Within their vehicles the ECUs — such as the Autopilot Control Unit, the Media Control Unit, and several Body Control Units — are interconnected through Ethernet and are much less than the usual within vehicles up to that point.
@@
The Volkswagen group take on SDV is based on creating and controlling the whole platform — cloud (Volkswagen Automotive Cloud VW.AC), hardware and software (VW. Operating System (OS)) [27], affecting interoperability with other automakers and hardware providers. While in initial stages, it was not possible to obtain documentation on APIs or Software Development Kits that should be used to developed vehicular services using the CARIAD ecosystem, highlighting a possible tendency towards a close proprietary SDV paradigm. A notable exception is the Roobadge [^13] Product description available at https://www.volkswagen.com.au/en/ roobadge.html., a solution to avoid collisions with kangaroos — very common in Australia — by adding, through a front badge or license plate holder sensors to detect the kangaroos. This solution not only showcases a real example of the potential of SDV, since new features are added to an existing vehicle through a new deployed service, but is also compatible with other brand’s vehicles, showcasing a real SDV example that considers brand interoperability.
@@
The Renault group considers a backbone of ECUs interconnected with Ethernet, with gateways to bridge the Ethernet connected ECUs with legacy sensors or actuators interconnected by CAN [19, 28]. These ECUs run Android and are based on Qualcomm chipsets, supporting an interoperable, open SDV architecture [^14] More details at www.qualcomm.com.. Volvo and Polestar have a similar approach, with zone controllers — high performance controllers in specific zones of the vehicle — interconnected and connected to a central master Vehicle Control Unit (VCU) with an Ethernet backbone [29]. The Zone Controllers can also act as a bridge to legacy buses, like CAN and Local Interconnect Network bus (LIN). Volvo within their SDV efforts also published OpenAPI specifications for their CAVs [30].
The BMW group also aims to go from signal oriented to unified vehicle wide layered Service Oriented Architecture (SOA) architecture with providers and consumers [24, 31]. This also requires latencies, configurations and bandwidth options encoded in the system.
The Stellantis group aims at supporting sustainability and user experience use cases through service oriented solutions. SOA can provide the necessary scalability previous service development paradigms could not — namely 1 application per ECU — while allowing a firm contract between services [32].
@@
General Motors is also working towards their SDV efforts through the uServices solution, a set of APIs for developers and Ultifi, a end-to-end software platform that “will enable Over-the-air (OTA) updates and in-car subscription services” [33].
@@
Non automakers are also working towards providing automakers with new tools. Nvidia is also entering the market with a clear AI first approach [^15] Developer documentation available at https://developer.nvidia.com/drive. — the Nvidia DRIVE OS, which includes a embedded Real-Time Operating System (RTOS), a hypervisor, Compute Unified Device Architecture (CUDA) libraries and TensorRT, and features such as secure boot, security services, firewall, and OTA updates [^16] Developer documentation available at https://developer.nvidia.com/drive/drive-sdk.. Vector [^17] Available at https://www.vector.com/ within its ecosystem propose tools to support development of vehicles with ECUs interconnected by Ethernet with TSN and of services using AUTOSAR within SOA architectures.
MotionWise [34] proposes a middleware for safe orchestration of vehicle software across ECUs for automated driving platforms, with support for AUTOSAR classic and adaptive, Robotic Operating System (ROS), and Data Distribution Service (DDS) support.
@@
The Eclipse Foundation, within its SDV working group, is also working within incubating initiatives to define orchestrators and virtualization managers (Ankaios, Velocitas), vehicle abstraction models (Chariott), a development platform (Kuksa), an abstraction layer for communication between nodes (eCAL) and operation system (MoEc.OS) with support for DDS and integration with environments such as the SUMO and CARLA simulators (further details in section VI-B). While in an initial development phase, these projects can be a good future basis for anyone entering the SDV.
Continental, with the Amazon Internet of Things (IoT) for Automotive [^18] More details at https://catalog.workshops.aws/awsiotforautomotive/en-US. support, is also working towards a “a modular multi-tenant hardware and software framework that connects the vehicle to the cloud” [^19] More details at https://aws.amazon.com/pt/blogs/architecture/developing-a-platform-for-
software-defined-vehicles-with-continental-automotive-edge-caedge/
Autoware, through its Open AD Kit, the first Scalable Open Architecture for Embedded Edge (SOAFEE) blueprint, enables cloud-native DevOps of autonomous driving solutions for SDV [^20] Developer documentation at https://autowarefoundation.github.io/open-ad-kit-docs/latest/version-2.0/..
Hardware companies such as Renesas propose the AosCloud [^21] More details at https://www.renesas.com/us/en/software-tool/aosedge-platform., a solution to support OEMs and service providers from seamless software installation to operation in the vehicle, and implement various new services and functional enhancements simply.

### v2

# Deterministic and Reliable Software-Defined Vehicles: key building blocks, challenges, and vision

@@ Abstract
As vehicle systems become increasingly complex, with more features, services, sensors, actuators, and processing units, it is important to view vehicles not just as modes of transportation moving toward full autonomy, but also as adaptive systems that respond to the needs of their occupants. Vehicular services can be developed to support these adaptations. However, the increasing complexity of vehicular service development, even with current standardizations, best practices and guidelines, are insufficient to tackle the high complexity of development, with expectations of up to 1 (U.S.) billion lines of code for a fully (level 5) autonomous vehicle.
Within this survey, the paradigm of Deterministic Software Defined Vehicles is explored, aiming to enhance the quality and ease of developing automotive services by focusing on service-oriented architectures, virtualization techniques, and the necessary deterministic intra- and inter-vehicular communications. Considering the main open challenges for such verticals, a vision architecture towards improved services development and orchestration is presented, focusing on: a) a deterministic network configurator; b) a data layer configurator; c) a hypervisor configurator; d) the vehicle abstraction layer; and e) a software orchestrator.
keywords:
Software-Defined Vehicles, Connected Vehicles, Vehicular Applications, Vehicular Services, Intelligent Transportation Systems, Deterministic Networks, Time Sensitive Networks, Survey
journal: Computer Communications
[…]
[overall]organization=Department of Electronics, Telecommunications and Informatics, University of Aveiro / Instituto de Telecomunicações Aveiro, addressline=Campus Universitário de Santiago, city=Aveiro, postcode=3810-193, country=Portugal
[…]
[duarte]organization=Instituto de Telecomunicações Aveiro, addressline=Campus Universitário de Santiago, city=Aveiro, postcode=3810-193, country=Portugal
@@
1 Introduction
[…]
@@ 1 Introduction
This growing code base is becoming unmanageable to vehicle manufacturers and developers of vehicular services within a SDV. As described in fig. 2, automakers will tend to leave unchecked, unverified and poorly tested software go into production vehicles. If the growth trend of the code base continues, the results can be potentially deadly [^3] Unverified logic and deadly effects on Tesla’s autopilot per washingtonpost.com.. The overall intra-vehicular systems, as developed following legacy standards and practices, are deemed too complex and static — and thus costly to maintain, update and keep safe — by vehicle manufacturers, with the reduction of the overall complexity (both in software and hardware terms, in particular wiring harness weight savings) becoming a major goal in order to reduce operational costs and decrease potential security issues from unchecked, poorly tested code [5].
Standardization of best practices and development guidelines for intra vehicular software gained traction during the early 2000s [6]. In addition, programming methodologies and guidelines [^4] Such as the ASIL within ISO 26262 or MISRA-C, a set of guidelines for safety-related systems, such as automotive vehicles. Details at synopsys.com, misra.org.uk and synopsys.com. are critical and cannot be disregarded. However, they are no longer enough to tackle such complexity of development.
The context of deterministic pieces of software that carry out something within the vehicle or its environment can provide a solution to this problem. Vehicular services, since they may be critical, shall be deterministic in a broad sense in order to ensure that the result is exactly what was intended in the exact moment it was intended. In this article, we propose that the quality and easiness of the development of services for automotive can be improved using the paradigm of Deterministic Software Defined Vehicles. The main contributions of this survey are as follows:
@@
    Demonstrate the suitability and relevance of the software defined vehicles for a sustainable development of vehicular services.
@@
The remaining of this document is organized into the following sections. Section 2 introduces the background concepts, defining vehicular services and their characteristics, and the concept of SDV and additional information on the industrial landscape. It also shows how automakers or suppliers are facing the SDV paradigm. Next, section 3 presents the main building blocks for a SDV based vehicle (and vehicular services) development, and section 4 discusses on the main challenges. Section 5 follows with the main use cases and requirements that can be drawn from the use cases. Section 6 considers those requirements for the proposal of an architecture and of an evaluation process to validate the work. This considers the required hardware and software for both simulation and real-world testing. Finally, conclusions are provided in section 7.
@@
2 Software Defined Vehicles
At the core of SDV there is the concept that features within a vehicle are mostly defined through software. This follows a trend where vehicles, that started with little to no features, then evolved to features defined by mechanical components and by the 1980s started to define the features through software. Figure 3 summarizes this evolution.
Figure 3: Evolution of vehicular features over time, including levels of SDV. Partially based on [8, 9].
The SDV paradigm main ideas are not new. Prior to the advent of SDVs, the automotive industry focus on software was mostly restricted to connectivity between vehicles and the user’s smartphones to provide basic insights into the state of the vehicle [10]. By 1998, [11] considered intelligent vehicles as essential to autonomous transportation, designing a proposal of a vehicle capable of communicating with a centralized management control system. Also in 1998, [12] proposes that smart vehicles are a junction of communication architectures, system architecture, sensors, tasks, kinematics and man-machine interface. The authors propose a communication architecture in a producer-distributor-consumer pattern with buses such as RS232 and Controller Area Network bus (CAN), where time coherence is necessary. By 2001, a reference model architecture for intelligent vehicles [13] considered the vehicle as having intelligent control systems with each layer with one or more computational nodes interconnected and connected to sensors.
@@ 2 Software Defined Vehicles
SDV aims to improve cars along their life-cycle [14], while potentially mitigating the increased unmanaged complexity of modern vehicular software, hardware, and communications. The paradigm states that vehicles manage their operations, add functionality, and enable new features primarily or entirely through software [^5] More details at blackberry.qnx.com.. Since vehicles evolve over their lifetime, with more features, solutions that may require remote control or remote updates, main features being delivered by software adds flexibility, reducing the difficulty of adding new features and allowing new business models to be considered by automakers during the lifetime of the vehicle.
The shift towards SDV means that new features can be pushed to existing and new vehicles, which are now more defined through software than through hardware or mechanical systems as in the past [3]. Not only this adds flexibility to vehicle manufacturers, but also creates new business models where owners can pay for features after buying the vehicle and during its lifetime. This follows a trend where vehicles are now seen by potential owners in a subscription payment model, instead of the one-time payments or leasing/pay-per-use models (fig. 4).
Figure 4: Preferred payments scheme to purchase/activate connectivity services, % of respondents (n = 1649) showcasing a shift in preferred ownership payment models [15].
@@
Such support not just improves the development of vehicular services, but also helps towards a decrease in hardware complexity (contributing towards a much needed decrease in the number of ECUs, cables and the cost and weight savings which follow), with more uniform software code-base (and therefore reduction of the costs and complexity of the development) and greater adaptability (essential specially if the future of the car ownership is car sharing [16]). This adaptability also supports a greater capability to process higher volumes of data while fulfilling the overall performance requirements — essential in Connected Autonomous Vehicles.
2.1 Industrial landscape
@@
Vehicular manufacturers are aware of the necessary changes in thinking. However, the transition from the previous mentality of closed source, customized solutions for each model of vehicle towards a more extendable, interoperable, architecture is far from being over. The Renault, Volvo/Polestar and Volkswagen groups seems to be ahead in terms of a more interoperable architecture based on ECUs interconnected through Ethernet that run Android and well-known versions of chipsets. It is however not possible to ignore efforts from other vehicle manufacturers that may be developing equivalent solutions but have not, at the time of this survey, announced their efforts. The opacity of the vehicular industry, while expected within a competitive market, it remains and will remain a threat to any related work, including this work. To mitigate this threat, research of gray literature is essential, targeting not only automakers but their suppliers and open source community initiatives.
Figure 5 summarizes the efforts of automakers on adoption the main aspects related to SDV, with table 1 providing further detail per associated areas.
@@
Table 1: Overview of the manufacturers effort on SDV.
@@
Existing automakers were not agile enough within the development of their vehicular services [17], which lead to the rise of new automakers, such as Tesla, in very software-intensive tasks such as autonomous driving. The Tesla company is in fact probably the most well-known example of application of the SDV paradigm principles. Within its vehicles, the ECUs — such as the Autopilot Control Unit, the Media Control Unit, and several Body Control Units — are interconnected through Ethernet, and are in a lower number than the usual within vehicles up to that point. This zonal architecture supports Advanced Driving Assistance System (ADAS) systems for level 2+ autonomous driving, such as the Full Self-Driving (Supervised) that is then sold to end-users in a subscription manner [^6] Details at tesla.com/support/full-self-driving-subscriptions..
As a standardized ”standardized software framework and open E/E system architecture”, the Automotive Open System Architecture (AUTOSAR)777autosar.org is a de facto standard, with the support from vehicle manufacturers and vendors. It is a highly complex ecosystem, with more than 100 modules and its documentation reaching more than 21 000 pages [6]. The standard has two different types of platforms as depicted in fig. 6: the Classic, for traditional hard real-time systems based on micro-controllers with a static configuration, and set of pre-compiled monolithic binaries; and also a newer Adaptive platform with greater flexibility and adaptability. The Classic platform is aimed at safety critical systems with hard real-time requirements with a stated linkage between services and CPU cores. Development through the original Classic Platform is however, very hard-coded to the vehicle and its Electrical/Electronic (E/E) architecture [18]. The Adaptive is capable of considering heterogeneous high performance ECUs and departs from legacy buses [19] such as CAN towards a Service Oriented Architecture (SOA) communication architecture (see  section 3.2) with Ethernet based on Scalable service-Oriented MiddlewarE over IP (SOME/IP) as bus protocol [20] and is overall more dynamic but less time predictable [21, 22], essential within critical vehicular systems.
Figure 6: AUTOSAR Classic VS Adaptive.
After the landscape changing forced by the United States through Tesla, European and Asiatic automakers such as BMW [23]; the Renault group through the Renault Software Factory [^8] More details on www.reuters.com.; Volvo; Kia and Hyundai [^9] More details on www.carscoops.com.; and the Volkswagen group through the CARIAD [^10] Available at cariad.technology., supported by companies such as Vector [24] and VXworks, are shifting towards a software-first development of the vehicles, based on novel hardware and communication paradigms.
@@
A major player within the European market, the Volkswagen group take on SDV is based on creating and controlling the whole platform — cloud (Volkswagen Automotive Cloud VW.AC), hardware and software (VW. Operating System (OS)) [25], affecting interoperability with other automakers and hardware providers. While in initial stages, it highlights a possible tendency towards a close proprietary SDV paradigm. However, it was not possible to obtain documentation on APIs or Software Development Kits that should be used to developed vehicular services using the CARIAD ecosystem. A notable exception is the Roobadge [^11] Product description at https://www.volkswagen.com.au/en/ roobadge.html., a solution to avoid collisions with kangaroos — very common in Australia — by adding, through a front badge or license plate holder sensors to detect the kangaroos. This solution not only showcases a real example of the potential of SDV, since new features are added to an existing vehicle through a new deployed service, but is also compatible with other brand’s vehicles, showcasing a real SDV example that considers brand interoperability (even if the actual implementation is closed).
@@
Still within Europe, the Renault group considers a backbone of ECUs interconnected with Ethernet, with gateways to bridge the Ethernet connected ECUs with legacy sensors or actuators interconnected by CAN [14, 26]. These ECUs run Android and are based on Qualcomm chipsets, supporting an interoperable, open SDV architecture [^12] More details at www.qualcomm.com.. Volvo and Polestar have a similar approach, with zone controllers (more in section 3.1) — high performance controllers in specific zones of the vehicle — interconnected and connected to a central master Vehicle Control Unit (VCU) with an Ethernet backbone [22]. The Zone Controllers can also act as a bridge to legacy buses, like CAN and Local Interconnect Network bus (LIN). Volvo within their SDV efforts also published OpenAPI specifications for their CAVs [27].
Other European players include Mercedes-Benz, BMW and Stellantis. Mercedes-Benz, after a transition to a zonal E/E architecture (with separations between infotainment, body and comfort, ADAS and interior domains), is working on the MB.OS, a proprietary chip-to-cloud platform capable of ensuring decoupling software from hardware. This solution is already allowing Over-the-air (OTA) updates to significantly change the MB User Experience (MBUX), with future focus on standardization efforts [28]. The BMW group also aims to go from signal oriented to unified vehicle wide layered SOA architecture with providers and consumers [23, 29]. This also requires latencies, configurations and bandwidth options encoded in the system. The Stellantis group aims at supporting sustainability and user experience use cases through service oriented solutions. SOA can provide the necessary scalability previous service development paradigms could not — namely 1 application per ECU — while allowing a firm contract between services [30], as described in section 3.2.
@@
Within the United States of America, General Motors is working towards their SDV efforts through the uServices solution, a set of APIs for developers and Ultifi, a end-to-end software platform that “will enable OTA updates and in-car subscription services” [31].
@@
When on the tier-2+ suppliers for automakers, it is clear they are also working towards providing automakers with new tools. Continental, with the Amazon Internet of Things (IoT) for Automotive [^13] More details at catalog.workshops.aws/awsiotforautomotive/. support, is also working towards a “a modular multi-tenant hardware and software framework that connects the vehicle to the cloud” [^14] More details at aws.amazon.com.. Nvidia is entering the market with a clear AI first approach — the Nvidia DRIVE OS, which includes a embedded Real-Time Operating System (RTOS), a hypervisor, Compute Unified Device Architecture (CUDA) libraries and TensorRT, and features such as secure boot, security services, firewall, and OTA updates [^15] Developer documentation at developer.nvidia.com/drive. Vector [^16] At vector.com within its ecosystem propose tools to support development of vehicles with ECUs interconnected by Ethernet with TSN and of services using AUTOSAR within SOA architectures. MotionWise [32] proposes a middleware for safe orchestration of vehicle software across ECUs for automated driving platforms, with support for AUTOSAR classic and adaptive, Robotic Operating System (ROS), and Data Distribution Service (DDS) support. Hardware companies such as Renesas propose the AosCloud [^17] More details at renesas.com., a solution to support Original Equipment Manufacturers and service providers from seamless software installation to operation in the vehicle, and implement various new services and functional enhancements simply.
@@
Regarding the open-source community, alternatives such as the Eclipse Foundation, within its SDV working group, is also working within incubating initiatives to define: a) orchestrators and virtualization managers (Ankaios, Velocitas); b) vehicle abstraction models (Chariott); c) a development platform (Kuksa); and d) an abstraction layer for communication between nodes (eCAL) and operating system (MoEc.OS) with support for DDS as well as simulators such as SUMO and CARLA (see section 6.2). While in an initial development phase, these projects can be a good future basis for anyone entering the SDV. The Scalable Open Architecture for Embedded Edge (SOAFEE)181818soafee.io, more details at arm.com. presents a cloud-native alternative with an architecture enhanced for mixed-criticality automotive applications with corresponding open-source reference implementations to enable commercial and non-commercial offerings. Autoware, an open-source project for autonomous driving191919autoware.org, through its Open AD Kit, the first SOAFEE blueprint, enables cloud-native DevOps of autonomous driving solutions for SDV [^20] Developer documentation at autowarefoundation.github.io..